3.Find the coordinate of the point of interception of the lines: x + 2y = 5 ; x + y = 3
a.(1,2)
b.(−1,2)
c.(1,1)
d.(1,0)

4.Two lines are said to be parallel if
a.their gradient is the same
b.they have equal rate of change
c.the angle between the two lines is zero
d.All the options
1
Expert's answer
2019-04-10T09:27:13-0400

3. To find the point of intersection, we need to solve the system of the equations of these two lines.

x + y = 3 -> y = 3 - x,

x + 2(3 - x) = 5,

x = 1, y = 3 - 1 = 2

The point is (1;2), so the correct answer is a.

4.Two lines are said to be parallel if their gradient is the same or they have equal rate of change and the angle between the two lines is zero, so the correct answer is d. All the options.